HUBBLE SPACE TELESCOPE DAILY REPORT #2874 PERIOD COVERED: 0000Z (UTC) 05/16/01 - 0000Z (UTC) 05/17/01 Daily Status Report as of 137/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.2 Completed STIS/CCD 8663 (Survey of SMC Planetary Nebulae: Nebular and Stellar Evolution in a Low- Metallicity Environment) The Space Telescope Imaging Spectrograph (CCD) was used to survey of SMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, in an environment that is chemically very metal poor. We will obtain STIS imaging and medium-resolution slitless spectroscopy that will yield line fluxes and nebular morphologies in important emission lines, plus magnitudes of the central stars. From these data we will gather a harvest of information: the nebular size, morphology, ionization structure, density, and mass; and the central star temperature, luminosity, and mass. The proposal completed with no reported anomalies. 1.3 Completed STIS/CCD 8700 (A STIS Spectroscopic Snapshot Survey of 3CR Radio Galaxies : The Nature of the Unresolved Nuclei) The Space Telescope Imaging Spectrograph (CCD) was used to perform STIS snapshot long-slit spectroscopy of these nuclei over 2900- 10300 AA to classify them {QSO, NLRG, BLRG, BL Lac, ldots} and establish their AGN characteristics down to the lowest nuclear luminosities, while still within powerful 3C radio galaxies. The primary goals are : {1} to test unification schemes by identifying previously undetected signatures of low-level AGN/QSO activity in their spectra in the form of weak, broad emission lines, featureless continua, and UV excess, {2} to constrain the FR I/BL Lac connection by measuring the relative contribution of thermal {starburst}, non-thermal {optical synchrotron}, and line emission, and {3} to perform line and continuum diagnostics to determine the dominant ionization mechanisms in the circumnuclear regions {tenths of arcseconds} and the role of dust extinction in the NLRGs and BLRGs in the form of geometrically thick or thin tori and foreground dust. The proposal completed as planned, with no reported anomalies. 1.4 Completed Five Sets of WF/PC-2 8824 (Decontaminations and Associated Observations Pt. 3/4) The WF/PC-2 was used for the monthly WFPC2 decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, UV throughput check, VISFLAT sweep, and internal UV flat check. There were no reported problems. 1.5 Completed Three Sets of WF/PC-2 8828 (Cycle 9 Supplemental Darks pt3/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.6 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to obtain UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. A little over halfway through this proposal, the STIS safing event occurred at 136/055215Z, affecting all observations after that point. 1.7 Completed FGS/1 9240 (Astrometric Mass Determination of an Extrasolar Planet Candidate) Fine Guidance Sensor-1R was used to perform astrometric observations that will allow a quick measure of the sky-plane motion of 55 Cancri, a G star with an Msin{i} = 0.9 MJ radial velocity companion with a 14 day period. By using 7 pairs of visits, where each visit within a pair has the same parallax factor, we will be able to accurately determine the stars's proper motion. The reflex motion of the star, due to its orbit about the star-companion barycenter, will manifest itself as a systematic residual to the measured proper motion if the companion is sufficiently massive. There were no reported problems. 1.8 Completed Two Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.9 Completed WF/PC-2 9249 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal designed to simulate scientific plans. The proposal completed with no reported problems. ********************************************************** The following scheduled proposals were not completed successfully due to the STIS safing event: STIS/CCD 8865 (Bias Monitor-Part 2) Two Sets of STIS/MA1 8643 (Ultraviolet Properties of the Metal Rich M87 Globular Cluster System) STIS/CCD/MA2 8681 (Snapshot Survey of Variability of Narrow and Broad Associated Absorption Lines in Quasars) Three Sets of STIS/CCD 8845 (Spectroscopic Flats C9) STIS/CCD 8687 (Elliptical Galaxies With Nuclear Disks of Stars: Black Hole Search and Stellar Populations) STIS/CCD 8262 (Nuclear Kinematics of the Dense Globular Cluster M-15) ********************************************************** 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 16 Successful: 16 2.3 Operations Notes: The Space Telescope Imaging Spectrograph (STIS) remains in safe. The safing event occurred at 136/055215Z. Analyses indicate that a fuse on the STIS power input line (spacecraft side) has failed. There is a redundant power path into the STIS. That path is active now, but only to supply heaters to keep the STIS warm. It is possible to switch all the equipment in the STIS to the redundant path. But, until it has been determined what may have caused the failure, the transfer of STIS to the redundant side will be delayed. Meetings to determine the safest method of returning STIS to an operational state are on-going, although it now appears likely that switching over to the redundant power feed may take several weeks in order to carry out the ground testing of the procedures, verify the flight software loads, and make other necessary changes. Schedulers at the STScI are rebuilding next week's observing calendar to include only WF/PC-2 observations. The STIS safing event took place during the proposal described in 1.6. Scheduled STIS proposals not completed successfully are listed at the bottom of those that did in Section 1. Using ROP SR-1A, the SSR EDAC error counter was cleared five times. HSTAR 8223 reports that the parameter OPSSCV2 flagged out-of-limits high at 136/214903Z and was back within normal ranges at 136/214933Z. Later, the parameter was out-of-limits high for the interval 136/215233Z to 136/215333Z. HSTAR 8224 documents an out-of-limits low condition for OOBBHT6T beginning at 136/221217Z, hitting 10.875 degrees. Operations personnel are monitoring. Further action will be taken should the parameter drop to 5 degrees. The NSSC-1 status buffer was dumped at 137/0001Z per ROP NS-03. Later, at 137/0324Z, the status buffer was dumped and cleared, using the same ROP. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW